OpenStudy (anonymous):

HELP: write an equation of the line that passes through the given point and has the given slope: (2,4) slope 3

OpenStudy (anonymous):

use the slope intercept form. y-y1 = m (x- x1)

OpenStudy (anonymous):

y-4= 3 ( x- 2)

OpenStudy (anonymous):

where m is the slope.

OpenStudy (anonymous):

try wolframalpha.com for problems like these. it gives the answer, plus walks you through the problem without having to wait. i love it

OpenStudy (anonymous):

so.. y- 4 = 3x - 6 or 3x-y-2=0